7. x?y+ xy– 9y = x3

Solve the Euler ODE
Answer #1

x²y + xy - gy= x² at x = ea - zaloge dy , dy dz - dy en dx pady dz Now (D 20-1) + D-9 ) y = 203 (p2 - D + p- 3) = 33 g)Р. y, (2) - - D 2-3 234 хез 2х3 Exez Чус) - I e (1-) x2 = + x3 logne Solution. уч) = 8 ) + ) У) = 4° с. х + - Prx